OpenStudy (babyslapmafro):

Please help me solve the following problem. Find two unit vectors that are normal to the given surface at the point P(3,5,1). Click to see the equation...

OpenStudy (babyslapmafro):

\[z^2=\sqrt{\frac{ z+x }{ y-1 }}\]

OpenStudy (dan815):

okay you can take the gradient of the surface at that point and normalize it

OpenStudy (dan815):